OpenStudy (anonymous):

What is the slope of this equation? y=2/5x

OpenStudy (saifoo.khan):

2/5

OpenStudy (saifoo.khan):

0.4

OpenStudy (anonymous):

Is that the answer

OpenStudy (saifoo.khan):

100%

OpenStudy (anonymous):

You Can use the Equation y=mx+b. Where M is the slope and b is the y value. Well, there is no B, so the equation here is y=mx, well M is filled in for 2/5, so 2/5 is your slope.

OpenStudy (anonymous):

Thanks that was helpful.!

OpenStudy (anonymous):

No problem :) anytime. If you need more help on geometry or algebra math, message me and ill be GLAD to help :) ok?

OpenStudy (anonymous):

Ok I will thanks for telling me that I think we will be talking more often lol

OpenStudy (anonymous):

Haha ok, here ill see if I tought you anything, or if you get it now. Here's a problem. And try not to look at the message above. Unless you need the help. On a graph, you realize that the slope is 4/5 and your y-intercept (y value) is 6. Write the equation in slope intercept form. HINT. Slope intercept form = Y=MX+B.

OpenStudy (anonymous):

Um...Are you serious you only gave me 1 answer. Mr Aussie is it??

OpenStudy (anonymous):

What do you mean one answer? I didnt mean it was multiple choice, I meant you write the answer. Well like i said, M is the slope, and in the problem the slope is 4/5, and B is the y-intercept. The y intercept is 6, so now that we know that we plug it in to the problem. (X is never included, it always remains x.) So The correct slope intercept form would be y= 4/5x + 6. Do you get it or do you need me to decribe it in more detail?

OpenStudy (anonymous):

Is it 6.8?

OpenStudy (anonymous):

Math is not my Smartest Subject.

OpenStudy (anonymous):

nope. You never solve for y. Or the slope, you leave it the same, once you find M, you do not divide. The slope is always 4/5.

OpenStudy (anonymous):

Well its not always 4/5, but in this equation it is. But the slope is always in fraction form. And its ok, its my strongest, so thats why im here to help :)

OpenStudy (anonymous):

Thanks your soo nice and Ok I'll try to remember that.

OpenStudy (anonymous):

aww *blushes* lol. but i like heping others, becaue one day, im gonna need the help and maybe someone would help me. Treat people the way you wish to be treated. But now, since i think you got it. Lets see: Write the equation in slope intercept form (y=mx+b) The y intercept if 4 and the slope is 1/9. What is the answer? In full equation form.

OpenStudy (saifoo.khan):

should i giv the ans?

OpenStudy (anonymous):

Nahh, not just yet, lets give her a chance to solve it and if she cant get it, then yes. :)

OpenStudy (saifoo.khan):

oh okay okay! =)